# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v3

@@ II-B Channels and channel sequences
We define a channel $W$ as a collection of conditional distributions $W(y|x)$, defined $\forall x\in{\mathcal{X}}$, $\forall y\in{\mathcal{Y}}$. Throughout the paper, we assume a finite set ${\mathcal{X}}$ and a countable set ${\mathcal{Y}}$. Formally, a channel $W$ is a function $W:{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, such that
$\forall x\in{\mathcal{X}}:\sum_{y\in{\mathcal{Y}}}W(y|x)=1.$
The fact that $W$ has input messages from ${\mathcal{X}}$ and output from ${\mathcal{Y}}$ is denoted by $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$.
Channel $W$ defines the conditional distribution of the output given its input. If we further define some input distribution $p\in{\mathbb{D}}_{\mathcal{X}}$, then we can compute the joint distribution of input and output, which is denoted by $p\circ W$:
@@
$p\circ W(*,y)=\sum_{x\in{\mathcal{X}}}p(x)W(y|x),$ (1b)
where (1b) defines the probability of output $y$, given the input distribution $p$.
@@ II-C Mutual information capacity
$\mathbf{i}(W,p,x,y)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{cases}0,\text{if }W(y|x)=0\text{ or }p\circ W(*,y)=0\\ \log_{2}\frac{W(y|x)}{p\circ W(*,y)},\text{ otherwise}\end{cases}$ (3)
where $p\circ W(*,y)$ is given by (1b). The mutual information of $W$ under input distribution $p$ is the expectation of mutual information density:
${\mathbf{I}}(W,p)$ $\stackrel{{\scriptstyle\triangle}}{{=}}{\mathbf{E}}_{p\circ W(x,y)}[\mathbf{i}(W,p,x,y)]=\sum_{x\in{\mathcal{X}},y\in{\mathcal{Y}}}p\!\circ\!W(x,y)\cdot\mathbf{i}(W,p,x,y).$ (4)
The mutual information capacity/i-capacity [4] of channel $W$ is the maximum possible mutual information of $W$ under any input distribution, and is given by [^3] In [4], the setup is for the much more general case, and “$\sup$” over the input distributions is used instead of “$\max$”. In our setup we assumed (see Section II-B) that ${\mathcal{X}}$ is finite and ${\mathcal{Y}}$ is countable, so the maximum is always achieved as a maximum of a continuous function over a compact set.
${\mathbf{I}}(W)\stackrel{{\scriptstyle\triangle}}{{=}}\max_{p\in{\mathbb{D}}_{\mathcal{X}}}{\mathbf{I}}(W,p).$ (5)
We will call the distribution $p$, for which the maximum is achieved, as the optimal distribution (for $W$). By the mutual information capacity/i-capacity of channel sequence $V=V^{\infty}$ we mean the asymptotic average i-capacity of $V_{n}$ per input symbol:
${\mathbf{I}}(V)\stackrel{{\scriptstyle\triangle}}{{=}}\lim_{n\to\infty}\frac{{\mathbf{I}}(V_{n})}{n},$ (6)
if the limit exists. Note that the i-capacity of a channel always exists, since ${\mathbf{I}}(W,p)$ is a concave function of the input distribution $p$ with countable number of coefficients; however, the i-capacity of a channel sequence might not exist, since sequence ${\mathbf{I}}(V_{n})/n$ might not converge [^4] Consider a channel which is ideal for even $n$ and complete noise for odd $n$. In this case, the sequence of values of ${\mathbf{I}}(V_{n})/n$ is $(\log_{2}|{\mathcal{X}}|,0,\log_{2}|{\mathcal{X}}|,0,\dots)$. Such sequence has no limit for $|{\mathcal{X}}|>1$..
@@ II-D Coding capacity
For a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$, code size $M\in{\mathbb{N}}$, $M\leq|{\mathcal{X}}|$ and error probability $\varepsilon\in[0,1]$, we define a $(W,M,\varepsilon)$-coding scheme as a collection $(x_{1}^{M},{\mathcal{R}}_{1}^{M}=({\mathcal{R}}_{1},{\mathcal{R}}_{2},\dots,{\mathcal{R}}_{M}))$, where:
@@
    $\forall i\in{\left\{{1,\dots,M}\right\}}:\sum_{y\notin{\mathcal{R}}_{i}}{W(y|x_{i})}\leq\varepsilon.$
Consider a channel sequence $V^{\infty}$. We call $R\in{\mathbb{R}}_{\geq 0}$ an achievable rate over $V$, if
@@
We denote by ${\mathbf{R}}(V)$ the set of achievable rates $R$ for channel sequence $V$. By the coding capacity/c-capacity ${\mathbf{C}}(W)$ of the channel sequence $V$ we mean the supremum of all achievable rates [5]:
${\mathbf{C}}(V)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\sup{\mathbf{R}}(V)$ (8)
The existence of a $(W,M,\varepsilon)$-coding scheme obviously implies the existence of a $(W,M^{\prime},\varepsilon^{\prime})$-coding scheme for any $M^{\prime}\leq M$ and $\varepsilon^{\prime}\geq\varepsilon$. So, if $R\in{\mathbf{R}}(V)$, then $R^{\prime}\in{\mathbf{R}}(V)$ for any rate $R^{\prime}<R$. Thus, the set ${\mathbf{R}}(V)$ is a connected set, i.e., there always exists $\alpha\in[0,\log_{2}|{\mathcal{X}}|]$, such that ${\mathbf{R}}(V)\in{\left\{{[0,\alpha],[0,\alpha)}\right\}}$. This implies that ${\mathbf{C}}(V)=\alpha$, i.e., any channel sequence has the c-capacity.
In [3], the coding capacity (c-capacity) ${\mathbf{C}}(V)$ is derived in terms of mutual information density:
${\mathbf{C}}(V)=\sup{\left\{{\alpha\left|\lim_{n\to\infty}~{}\min_{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\Pr_{p_{n}\circ V_{n}(x,y)}{\left\{{\frac{\mathbf{i}(V_{n},p_{n},x,y)}{n}\leq\alpha}\right\}}=0\right.}\right\}}$ (9)
@@ II-E Information stability
A channel sequence $V$ is called $J$-information stable [2], where $J=J^{\infty}$ with $J_{n}\to\infty$, if there exist distributions $p^{*}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$ and sequence $\delta$ with $\delta_{n}>0$, $\delta_{n}\to 0$, such that
@@
If $J_{n}={\mathbf{I}}(V_{n})$ then $V$ is called information-stable. We denote by ${\mathbf{I}}{\mathbf{S}}(V)$ the fact that $V$ is information-stable, and, moreover, the i-capacity ${\mathbf{I}}(V)$ exists. In the case of ${\mathbf{I}}{\mathbf{S}}(V)$, under our general assumption that ${\mathcal{X}}$ is finite, (10) can be rewritten as
$\forall\delta>0:\lim_{n\to\infty}~{}\Pr_{p^{*}_{n}\circ V_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]}=0.$ (11)
@@
${\mathbf{I}}{\mathbf{S}}(V)\iff{\mathbf{C}}(V)={\mathbf{I}}(V)$.
For further derivations we need the following sufficient condition of information stability.
@@
${\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}>\delta\cdot\Pr{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]},$
so
$0\leq\lim_{n\to\infty}~{}\Pr{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]}\leq\frac{1}{\delta}\cdot\lim_{n\to\infty}~{}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0,$

### v5

@@ II-B Channels and Channel Sequences
Recall (see the beginning of Section II) that a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ is a collection of conditional distributions $W(y|x)$, defined for all $x\in{\mathcal{X}}$, $y\in{\mathcal{Y}}$. Throughout the paper, we assume finite ${\mathcal{X}}$ and countable ${\mathcal{Y}}$.
If we further define some input distribution $p\in{\mathbb{D}}_{\mathcal{X}}$, then we can compute the joint distribution of input and output and the distribution of output:
@@
$p\!\bullet\!W(y)$ $=\sum_{x\in{\mathcal{X}}}p\!\circ\!W(x,y).$ (7b)
@@ II-C Mutual Information Capacity
$\mathbf{i}(W,p,x,y)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{cases}0,\text{if }W(y|x)=0\text{ or }p\!\bullet\!W(y)=0\\ \log_{2}\displaystyle\frac{W(y|x)}{p\!\bullet\!W(y)},\text{ otherwise}\end{cases}$ (9)
The mutual information of $W$ under input distribution $p$ is the expectation of mutual information density:
${\mathbf{I}}(W,p)\;{\overset{\triangle}{=}}\;$ ${\mathbf{E}}_{p\circ W(x,y)}[\mathbf{i}(W,p,x,y)]$
[…]
The mutual information capacity/i-capacity [27] of channel $W$ is the maximum possible mutual information of $W$ under any input distribution, and is given by
${\mathbf{I}}(W){\overset{\triangle}{=}}\max_{p\in{\mathbb{D}}_{\mathcal{X}}}{\mathbf{I}}(W,p).$ (11)
We will call the distribution $p$, for which the maximum is achieved, the optimal distribution (for $W$). By the mutual information capacity/i-capacity of channel sequence $V=V^{\infty}$ we mean the asymptotic average i-capacity of $V_{n}$ per input symbol:
${\mathbf{I}}(V){\overset{\triangle}{=}}\lim_{n\to\infty}\frac{{\mathbf{I}}(V_{n})}{n},$ (12)
if the limit exists. Note that the i-capacity of a channel always exists, since ${\mathbf{I}}(W,p)$ is a concave function of the input distribution $p$ with countable number of coefficients; however, the i-capacity of a channel sequence might not exist, since sequence ${\mathbf{I}}(V_{n})/n$ might not converge [^2] Consider a channel which is ideal for even $n$ and complete noise for odd $n$. In this case, the sequence of values of ${\mathbf{I}}(V_{n})/n$ is $\log_{2}|{\mathcal{X}}|$, 00, $\log_{2}|{\mathcal{X}}|$, $0,\dots$ Such sequence has no limit for $|{\mathcal{X}}|>1$..
For a channel $W$ and its input distribution $p$, define the mutual information variance as the variance of mutual information density:
${\mathbf{V}}(W,p)=\underset{p\circ W(x,y)}{{\mathbf{E}}}{\left[{{\left({{\mathbf{i}}(W,p,x,y)-{\mathbf{I}}(W,p)}\right)}^{2}}\right]}.$ (13)
@@ II-D Coding Capacity
For a channel $W:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$, code size $M\in{\mathbb{N}}$, $M\leq|{\mathcal{X}}|$ and error probability $\varepsilon\in[0,1]$, we define a $(W,M,\varepsilon)$-coding scheme as a collection $(x_{1}^{M},{\mathcal{R}}_{1}^{M})$, where:
@@
    $\forall i\in{\left\{{1,\dots,M}\right\}}:\sum_{y\notin{\mathcal{R}}_{i}}{W(y|x_{i})}\leq\varepsilon.$ (14)
Consider a channel sequence $V^{\infty}$. Define the set ${\mathbf{R}}(V)$ of all achievable rates over $V$ as the set of $R\in{\mathbb{R}}_{\geq 0}$, such that
@@
By coding capacity/c-capacity ${\mathbf{C}}(W)$ of the channel sequence $V$ we mean the supremum of achievable rates [9]:
${\mathbf{C}}(V)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\sup{\mathbf{R}}(V),$ (16)
The existence of a $(W,M,\varepsilon)$-coding scheme obviously implies the existence of a $(W,M^{\prime},\varepsilon^{\prime})$-coding scheme for any $M^{\prime}\leq M$ and $\varepsilon^{\prime}\geq\varepsilon$. So, if $R\in{\mathbf{R}}(V)$, then $R^{\prime}\in{\mathbf{R}}(V)$ for any $R^{\prime}<R$. Thus, the set ${\mathbf{R}}(V)$ is a connected set, i.e., ${\mathbf{R}}(V)\in{\left\{{[0,\alpha],[0,\alpha)}\right\}}$ for some $\alpha$, and therefore any channel sequence has its coding capacity.
In [26], the coding capacity ${\mathbf{C}}(V)$ is derived in terms of mutual information density: ${\mathbf{C}}(V)$ is the supremum of all $\alpha$, for which
$\lim_{n\to\infty}~{}\min_{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\Pr_{p_{n}\circ V_{n}(x,y)}{\left\{{\frac{\mathbf{i}(V_{n},p_{n},x,y)}{n}\leq\alpha}\right\}}=0.$ (17)
@@ II-E Information Stability
A channel sequence $V=V^{\infty}$ is called $J$-information stable [25], where $J=J^{\infty}$ with $J_{n}\to\infty$, if there exist distributions $p^{*}_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$ and sequence $\delta^{\infty}$ with $\delta_{n}>0$, $\delta_{n}\to 0$, such that
@@
If $J_{n}={\mathbf{I}}(V_{n})$ then $V$ is called information-stable. If, furthermore, the limit ${\mathbf{I}}(V)$ of ${\mathbf{I}}(V_{n})/n$ exists, then we call $V$ a strictly information-stable channel sequence and write ${\mathbf{I}}{\mathbf{S}}(V)$ to denote this fact. In the case of ${\mathbf{I}}{\mathbf{S}}(V)$, under our general assumption that ${\mathcal{X}}$ is finite, (18) is equivalent to the following: for all $\delta>0$,
$\lim_{n\to\infty}~{}\Pr_{p^{*}_{n}\circ V_{n}(x,y)}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]}=0.$ (19)
@@
${\mathbf{I}}{\mathbf{S}}(V)$ iff ${\mathbf{C}}(V)={\mathbf{I}}(V)$.
For further derivations, we need the following sufficient condition of information stability.
@@
$0\leq\lim_{n\to\infty}~{}\Pr{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right]}$
$\leq\frac{1}{\delta}\cdot\lim_{n\to\infty}~{}{\mathbf{E}}{\left[{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(V)}\right|}\right]}=0,$ (21)